What is the resistivity of a cylindrical material wire whose 1.0m length has a resistance of 2.0 ohms?the diameter of the wire is 0.5ohms.
1
Expert's answer
2018-08-22T11:06:08-0400
ρ=(R·S)/l=(2 Ohms·0.5 〖mm〗^2)/(1 m)=1 (Ohms·〖mm〗^2)/m